OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 2x + 5 A. (5, 10) and (2, 9) B. (5, 15) and (9, 2) C. (15, 5) and (2, 9) D. (5, 15) and (2, 9)

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Hey there!

OpenStudy (anonymous):

The best way to find the solutions to the equation is to plug in the points in the answer choices!

OpenStudy (skullpatrol):

What does a solution mean to you @plaguex?

OpenStudy (anonymous):

an answer

OpenStudy (skullpatrol):

Yes, but the answer must be true, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I'll do one for you: (5,10) That is in the form of (x,y) so substitute x = 5, and y = 10 10 = 2(5) + 5 10 = 10 + 5 10 ≠ 15 So that means (5,10) is NOT a solution to y = 2x + 5. We can then go ahead and cross out all the answer choices that have (5,10) in it!

OpenStudy (skullpatrol):

Ok, substitute the values into the given equation and find the one that gives a true statement

OpenStudy (anonymous):

Go ahead and see if (2, 9) is a solution to the equation. :)